# Benchmarking LLMs’ Judgments with No Gold Standard

### v1

@@ 1. Introduction
Most benchmarks are based on multiple-choice questions or other questions with objective gold standard answers, since it is easy to verify the LLMs’ outputs. While they provide valuable evaluation for LLMs, open-ended tasks, e.g. providing judgment about a research paper, encompass a broader array of skills and require both objective and subjective reasoning. In addition, concern has been raised about data contamination (also called data leakage), where the training data contains information about the tasks in the benchmarks, as the LLMs are pre-trained on massive online data, which is also the source of some of the benchmark tasks. LLMs can show unreliably good performance due to data contamination (Sainz et al., 2023; Golchin and Surdeanu, 2023; Oren et al., 2023). In contrast, with open-ended questions, LLMs can be asked to provide judgments about newly created content, e.g. the latest academic papers, for which LLMs have yet to index human evaluations or responses.
@@
A straightforward method may be to use another LLM as an oracle examiner directly to provide the evaluation, which has shown effectiveness in evaluating open-response QA (Bai et al., 2024), chatbot (Zheng et al., 2023b), etc. However, LLM examiners are susceptible to certain manipulations. In our experiment, elongating all responses by adding the same fixed sentences can significantly increase the GPT-4o LM examiner’s score. Interestingly, previous research has shown that even human evaluations can be manipulated by adding meaningless text (Goldberg et al., 2023). Meanwhile, other automated natural language generation (NLG) evaluation metrics, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2023), rely on comparison with a gold standard reference.
1.1. Our Contributions
We propose the Generative Estimator for Mutual information (GEM) which uses the estimated Shannon mutual information (MI) between a set of candidate responses and a set of peer reference responses, which need not be of gold standard quality. The mutual information measures the amount of information revealed about the reference responses by obtaining the candidate responses. By artificially removing stylistic and syntactical information, our approach measures how much semantic information the candidate responses can reveal about the reference responses, which is related to the concept of semantic coverage (Yuan et al., 2021; Nenkova and Passonneau, 2004). We additionally propose a variant of our method, Generative Estimator for Mutual Information with Synopsis (GEM-S), which estimates the mutual information conditional on a synopsis of the task, e.g. the abstract of the paper. This prevents a candidate response from receiving a high score based solely on superficial information. Consequently, the score emphasizes the additional semantic information gained from the candidate responses.
@@ 1.1. Our Contributions
Our work creates a bridge between the information theoretical framework in the literature of information elicitation without verification (Kong and Schoenebeck, 2019; Lu et al., 2024) and the NLG evaluation problem. Though GEM and GEM-S resemble the GPPM and GSPPM mechanisms in Lu et al. (2024), with manipulation resistance aligned to their incentive compatibility, we make a necessary change to make the score more suitable for the NLG evaluation problem. Specifically, for incentive compatibility, the score can rely solely on accuracy of predicting the peer reference, while we use mutual information to capture the gain in that accuracy for evaluation purposes.
@@ 1.2. Related Work
Given the success of LLMs, they have the potential to assist with peer reviews when used appropriately (Robertson, 2023; Kuznetsov et al., 2024). Liang et al. (2023) use a survey to evaluate GPT-4-generated reviews, and find that the GPT-generated reviews are thought helpful by more than 50% of participants. However, there have been concerns regarding issues such as hallucinations (Donker, 2023) and inconsistent performance (Liu and Shah, 2023). While our study uses the peer review scenario to evaluate LLM capabilities, it also provides a quantitative comparison for the effectiveness of various LLMs in generating informative reviews.
@@
Both information elicitation and machine learning employ a principal-agent model (Ali and Silvey, 1966). Specifically, the principal aims to elicit information from the agent, e.g. the probability of rain tomorrow. To incentivize the agent to provide truthful and informative reports, payment mechanisms are employed that reward truthful and informative reports more than untruthful or uninformative ones.
When the information is verifiable, e.g., we will eventually know whether it rains tomorrow, proper scoring rules (Good, 1952; Brier, 1950; Hendrickson and Buehler, 1971) can be applied, which are similar to the loss functions in supervised learning. When the information is subjective and unverifiable, Miller et al. (2005) propose the peer prediction mechanism, suggesting rewarding data by how well it predicts a peer’s subjective report which is judged by a proper scoring rule. Prelec (2004) propose the Bayesian Truth Serum mechanism. Kong and Schoenebeck (2019) propose an information-theoretic framework, suggesting paying the agents according to the f-mutual information between their reports. The framework provides a unified theoretical view of previous mechanisms in Dasgupta and Ghosh (2013), and Prelec (2004). Lu et al. (2024) first generalize the peer prediction to elicit informative textual responses with LLMs, providing the foundation for our GEM metrics.
Most information elicitation mechanisms theoretically ensure that truthful and informative reports yield higher expected scores. However, high ex-post correlation between scores and report qualities is critical for practical application, as highlighted by Burrell and Schoenebeck (2021); Xu et al. (2024). In our study, to validate GEM’s effectiveness in evaluating textual reports, we measure both the ex-post correlation with human-annotated quality scores and expected score change after several degradation and manipulation strategies.
@@
Kong and Schoenebeck (2018b) bridges information elicitation without verification and co-training, by formulating an optimization problem that maximizes the mutual information between two Bayesian predictors with two data sources. Cao et al. (2019); Xu et al. (2019) apply this information theoretical approach in co-training and de-noising respectively. Concurrently and independently, Belghazi et al. (2018) and Hjelm et al. (2018) propose neural network-based methods for estimating mutual information, which can be used as an optimization goal in unsupervised and self-supervised learning settings.
2. Preliminaries
@@
To estimate the MI with samples, we employ Point-wise Mutual Information (PMI) (Fano and Hawkins, 1961; Church and Hanks, 1990), an unbiased estimator of the MI. Formally, the PMI between realization $x$ and $y$ is defined as
@@
3. Generative Estimator for Mutual Information (GEM)
@@
3.1. Theoretical Guarantees: Benchmarking LLMs with No Gold Standard
We now show that even when the reference is not gold standard, the estimated MI can still be a benchmark where better candidates achieve higher scores theoretically. We adopt a widely-used model in decision-making theory (Blackwell, 1951, 1953). The candidate’s response follows an information structure, a mapping $\sigma:\mathcal{W}\rightarrow\Delta\Sigma$, where $\Sigma$ is the response space. It represents the distribution of responses conditional on the task.
Within this model, Blackwell’s order (Blackwell, 1951, 1953) provides a partial order of the informativeness of information structures. Consider two candidates, H and L, whose responses $X_{H}$ and $X_{L}$ follow information structures $\sigma_{H}$ and $\sigma_{L}$ respectively. Information structure $\sigma_{H}$ is more informative than $\sigma_{L}$ in the sense of Blackwell order if there exists a stochastic mapping $\Gamma$, such that $\sigma_{L}=\Gamma\sigma_{H}$. Intuitively, the information structure with a lower Blackwell order is more noisy, and thus, can be regarded as lower quality.
@@ 3.1. Theoretical Guarantees: Benchmarking LLMs with No Gold Standard
The proof is provided in Appendix B, following previous information elicitation literature (Lu et al., 2024). It mainly relies on the data processing inequality (Shannon, 1948; Thomas and Joy, 2006). Additionally, we provide a similar proposition for conditional MI in Appendix B.
@@ 3.2. Filtering Out Shortcuts
Figure 2. Example of Hierarchical Information Structure in Peer Reviews
@@
Figure 3. An Overview of Our Generative Estimator for Mutual Information
@@
Kong and Schoenebeck (2018a) suggest surface-level information can inflate mutual information through simple correlations, reducing the mutual information’s effectiveness in reflecting semantic informativeness, especially when the responses have a hierarchical information structure as shown in Figure 2. As some superficial information can be contained in a synopsis of the task, e.g. the abstract of the paper to be reviewed, conditional mutual information can be used to distill the semantic informativeness (Lu et al., 2024).
@@
4. Experiment: Validating GEM’s Effectiveness
In this section, we first provide an overview of our experiment setup to validate the effectiveness of the GEM metric in benchmarking LLMs without gold-standard references, and show the results. We provide our code repository at https://github.com/yx-lu/Benchmarking-LLMs--Judgments-with-No-Gold-Standard.
@@ 4. Experiment: Validating GEM’s Effectiveness
For text preprocessing, we employ GPT-4o. For GEM-S, without further indication, we use the abstract of the paper as the synopsis in this section. The prompts for probability estimation and preprocessing can be found in Appendix C.
@@
4.1. Result 1: Positive Correlation with Human Annotation
@@
Score
For task $w_{i}$ with corresponding responses $\mathbf{x}_{i}=\{x_{ij}\}$ where $x_{ij}$ is the $j$-th response. The score of response $x_{ij}$ is computed by $\frac{1}{|\mathbf{x}_{i}|-1}\sum_{k\neq j}f(w_{i},x_{ij},x_{ik})$ where $f$ is the evaluation metric.
Human-Annotated Peer Grading Dataset
This dataset contains 30 student project proposals in a graduate-level class on machine learning and about 180 peer evaluations of the proposals. We use GPT-4o to generate a short abstract for each proposal. Each proposal has at least 4 reviews, and each review contains an overall score and textual feedback, including “Strengths of the project”, “Weaknesses of the project/likely roadblocks”, and “Ideas for improvement or specific directions”. According to the quality of students’ textual feedback, an instructor manually grades the reviews with grade A, B, or C.
@@
Table 1. Spearman’s correlation coefficient ($\rho$) and p-values between various evaluation metrics and instructor-annotated grades. Significant positive correlations ($p<0.05$) are bolded.
Statistics Metric and Results
Table 1 shows Spearman’s correlation coefficient $\rho$ between the scores of various evaluation metrics and the instructors’ grades. All variants of our GEM metrics show a significant positive correlation. The GEM-S has the highest correlation over all three GEM variants, implying the effectiveness of filtering out shortcuts. It also shows competitive performance compared to GPT-4o LMExaminer which has the highest correlation. All other baselines perform worse than GEM, and only BARTScore-Recall has a significant positive correlation. In later sections, we provide further comparisons between GEM and LMExaminer.  [^3] An intriguing future direction is to re-calculate the GEM score with rephrased system prompts or different preprocessing methods, and take the average to reduce the noise of GEM metric.
4.2. Result 2: Better Sensitivity to Degradation
In this section, we aim to compare how sensitive the evaluation metrics are to degradations that obviously reduce the semantic informativeness of the responses. A statistically significant score decrease indicates the sensitivity of the metric. We first introduce our validation workflow in Algorithm 1, the ICLR peer review dataset, and then introduce several degradation strategies. The workflow and dataset will also be used in the robustness check again manipulation strategies in Section 4.3.
@@ 4.2. Result 2: Better Sensitivity to Degradation
ICLR Dataset
We use the ICLR 2023444At the time of our experiment, ICLR 2024 review data was not yet available through the OpenReview API. peer review data publicly available on OpenReview. We randomly select 300 papers as our benchmark dataset, and for each paper, we randomly select 3 original human reviews: one review to serve as a human candidate, and the other two as reference responses.
Degradation Strategies
We introduce three degradation strategies as follows. Note that the degraded responses after Sentence Deletion or Deletion & Completion can be regarded as Blackwell-dominated by the original one. The detailed setup is provided in Appendix D.1.
  1. (1)
@@
Statistics and Results
We compare the standardized mean differences (SMD) of evaluation metrics in Tabel 2: GEM and GEM-S are only two metrics that consistently have significant negative SMD on all three degradations. The LMExaminer shows the highest effectiveness on Sentence Deletion and Deletion & Completion, but failed to penalize reviews only based on the abstract.
@@
Table 2. Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations and manipulations of various evaluation metrics. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation. Significant score increases ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric is not robust against the manipulation. We provide 95% confidence interval of SMD values in Appendix A.1.
4.3. Result 3: Better Robustness against Manipulation

### v2

@@ 1 Introduction
Most benchmarks are based on multiple-choice questions or other questions with objective gold standard answers, since it is easy to verify the LLMs’ outputs. While they provide valuable evaluation for LLMs, open-ended tasks, e.g. providing judgment about a research paper, encompass a broader array of skills and require both objective and subjective reasoning. In addition, concern has been raised about data contamination (also called data leakage), where the training data contains information about the tasks in the benchmarks, as the LLMs are pre-trained on massive online data, which is also the source of some of the benchmark tasks. LLMs can show unreliably good performance due to data contamination (Sainz et al., 2023; Golchin & Surdeanu, 2023; Oren et al., 2023). In contrast, with open-ended questions, LLMs can be asked to provide judgments about newly created content, e.g. the latest academic papers, for which LLMs have yet to index human evaluations or responses.
@@
A straightforward approach may involve utilizing an alternative LLM as an oracle examiner to directly provide the evaluation, which has demonstrated efficacy in evaluating open-response QA (Bai et al., 2024), chatbot (Zheng et al., 2023b), etc. However, LLM examiners are susceptible to certain manipulations. In our experiments, elongating all responses by adding the same fixed sentences can significantly increase the GPT-4o LM examiner’s score. Interestingly, previous research has shown that even human evaluations can be manipulated by adding meaningless text (Goldberg et al., 2023). Meanwhile, other automated natural language generation (NLG) evaluation metrics, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), BARTScore (Yuan et al., 2021) and GPTScore (Fu et al., 2023), rely on comparison with a gold standard reference.
1.1 Our Contributions
We propose the Generative Estimator for Mutual information (GEM) which uses the estimated Shannon mutual information (MI) between a set of candidate responses and a set of peer reference responses, which need not be of gold standard quality. The mutual information measures the amount of information revealed about the reference responses by obtaining the candidate responses. By artificially removing stylistic and syntactical information, our approach measures how much semantic information the candidate responses can reveal about the reference responses, which is related to the concept of semantic coverage (Yuan et al., 2021; Nenkova & Passonneau, 2004). We additionally propose a variant of our method, Generative Estimator for Mutual Information with Synopsis (GEM-S), which estimates the mutual information conditional on a synopsis of the task, e.g. the abstract of the paper. This prevents a candidate response from receiving a high score based solely on superficial information. Consequently, the score emphasizes the additional semantic information gained from the candidate responses.
@@ 1.1 Our Contributions
Our work establishes a bridge between the information theoretical framework in the literature of information elicitation without verification (Kong & Schoenebeck, 2019; Lu et al., 2024) and the NLG evaluation problem. Though GEM and GEM-S resemble the GPPM and GSPPM mechanisms in Lu et al. (2024), with manipulation resistance aligned to their incentive compatibility, we make a necessary change to make the score more suitable for the NLG evaluation problem. Specifically, for incentive compatibility, the score can rely solely on accuracy of predicting the peer reference, while we use mutual information to capture the gain in that accuracy for evaluation purposes.
@@ 1.2 Related Work
Given the success of LLMs, they have the potential to assist with peer reviews when used appropriately (Robertson, 2023; Kuznetsov et al., 2024). Liang et al. (2023) use a survey to evaluate GPT-4-generated reviews, and find that the GPT-generated reviews are thought helpful by more than 50% of participants. However, there have been concerns regarding issues such as hallucinations (Donker, 2023) and inconsistent performance Liu & Shah (2023). While our study uses the peer review scenario to evaluate LLM capabilities, it also provides a quantitative comparison for the effectiveness of various LLMs in generating informative reviews.
@@
Both information elicitation and machine learning employ a principal-agent model (Ali & Silvey, 1966). Specifically, the principal aims to elicit information from the agent, e.g. the probability of rain tomorrow. To incentivize the agent to provide truthful and informative reports, payment mechanisms are employed that reward truthful and informative reports more than untruthful or uninformative ones.
When the information is verifiable, e.g., we will eventually know whether it rains tomorrow, proper scoring rules (Good, 1952; Brier, 1950; Hendrickson & Buehler, 1971) can be applied, which are similar to the loss functions in supervised learning. When the information is subjective and unverifiable, Miller et al. (2005) propose the peer prediction mechanism, suggesting rewarding data by how well it predicts a peer’s subjective report which is judged by a proper scoring rule. Prelec (2004) propose the Bayesian Truth Serum mechanism. Kong & Schoenebeck (2019) propose an information-theoretic framework, suggesting paying the agents according to the f-mutual information between their reports. The framework provides a unified theoretical view of previous mechanisms in Dasgupta & Ghosh (2013), and Prelec (2004). Lu et al. (2024) first generalize the peer prediction to elicit informative textual responses with LLMs, providing the foundation for our GEM metrics.
Most information elicitation mechanisms theoretically ensure that truthful and informative reports yield higher expected scores. However, high ex-post correlation between scores and report qualities is critical for practical application, as highlighted by Burrell & Schoenebeck (2021); Xu et al. (2024). In our study, to validate GEM’s effectiveness in evaluating textual reports, we measure both the ex-post correlation with human-annotated quality scores and expected score change after several degradation and manipulation strategies.
@@
Kong & Schoenebeck (2018b) bridges information elicitation without verification and co-training, by formulating an optimization problem that maximizes the mutual information between two Bayesian predictors with two data sources. Cao et al. (2019); Xu et al. (2019) apply this information theoretical approach in co-training and de-noising respectively. Concurrently and independently, Belghazi et al. (2018) and Hjelm et al. (2018) propose neural network-based methods for estimating mutual information, which can be used as an optimization goal in unsupervised and self-supervised learning settings.
2 Preliminaries
@@
To estimate the MI with samples, we employ Point-wise Mutual Information (PMI) (Fano & Hawkins, 1961; Church & Hanks, 1990), an unbiased estimator of the MI. Formally, the PMI between realization $x$ and $y$ is defined as
@@
3 Generative Estimator for Mutual Information (GEM)
@@
3.1 Theoretical Guarantees: Benchmarking LLMs with No Gold Standard
We now show that even when the reference is not gold standard, the estimated MI can still be a benchmark where better candidates achieve higher scores theoretically. We adopt a widely-used model in decision-making theory (Blackwell, 1951; 1953). The candidate’s response follows an information structure, a mapping $\sigma:\mathcal{W}\rightarrow\Delta\Sigma$, where $\Sigma$ is the response space. It represents the distribution of responses conditional on the task.
Within this model, Blackwell’s order (Blackwell, 1951; 1953) provides a partial order of the informativeness of information structures. Consider two candidates, H and L, whose responses $X_{H}$ and $X_{L}$ follow information structures $\sigma_{H}$ and $\sigma_{L}$ respectively. Information structure $\sigma_{H}$ is more informative than $\sigma_{L}$ in the sense of Blackwell order if there exists a stochastic mapping $\Gamma$, such that $\sigma_{L}=\Gamma\sigma_{H}$. Intuitively, the information structure with a lower Blackwell order is more noisy, and thus, can be regarded as lower quality.
@@ 3.1 Theoretical Guarantees: Benchmarking LLMs with No Gold Standard
The proof is provided in Appendix B, following previous information elicitation literature (Lu et al., 2024). It mainly relies on the data processing inequality (Shannon, 1948; Thomas & Joy, 2006). Additionally, we provide a similar proposition for conditional MI in Appendix B.
@@ 3.2 Filtering Out Shortcuts
Figure 2: Example of Hierarchical Information Structure in Peer Reviews
@@
Figure 3: An Overview of Our Generative Estimator for Mutual Information
@@
Kong & Schoenebeck (2018a) suggest surface-level information can inflate mutual information through simple correlations, reducing the mutual information’s effectiveness in reflecting semantic informativeness, especially when the responses have a hierarchical information structure as shown in Figure 2. As some superficial information can be contained in a synopsis of the task, e.g. the abstract of the paper to be reviewed, conditional mutual information can be used to distill the semantic informativeness (Lu et al., 2024).
@@
4 Experiment: Validating GEM’s Effectiveness
In this section, we provide our experiment setup to validate the effectiveness of the GEM metric in benchmarking LLMs without gold-standard references, and show the results. Code repository is at https://github.com/yx-lu/Benchmarking-LLMs--Judgments-with-No-Gold-Standard.
@@ 4 Experiment: Validating GEM’s Effectiveness
For text preprocessing, we employ GPT-4o. We also present the results with Llama-3.2 90B preprocessing in Appendix A.3, and demonstrate the robustness of GEM metrics with various preprocessing models. For GEM-S, without further indication, we use the abstract of the paper as the synopsis in this section. The prompts for probability estimation and preprocessing can be found in Appendix C.
@@
4.1 Result 1: Positive Correlation with Human Annotation
@@
Score. For task $w_{i}$ with corresponding responses $\mathbf{x}_{i}=\{x_{ij}\}$ where $x_{ij}$ is the $j$-th response. The score of response $x_{ij}$ is computed by $\frac{1}{|\mathbf{x}_{i}|-1}\sum_{k\neq j}f(w_{i},x_{ij},x_{ik})$ where $f$ is the evaluation metric.
Human-Annotated Peer Grading Dataset. This dataset contains 30 student project proposals in a graduate-level class on machine learning and about 180 peer evaluations of the proposals. We use GPT-4o to generate a short abstract for each proposal. Each proposal has at least 4 reviews, and each review contains an overall score and textual feedback, including “Strengths of the project”, “Weaknesses of the project/likely roadblocks”, and “Ideas for improvement or specific directions”. According to the quality of students’ textual feedback, an instructor manually grades the reviews with grade A, B, or C.
@@
Table 1: Spearman’s correlation coefficient ($\rho$) and p-values between various evaluation metrics and instructor-annotated grades. Significant positive correlations ($p<0.05$) are bolded.
Statistics Metric and Results. Table 1 shows Spearman’s correlation coefficient $\rho$ between the scores of various evaluation metrics and the instructors’ grades. All variants of our GEM metrics show a significant positive correlation. The GEM-S has the highest correlation over all three GEM variants, implying the effectiveness of filtering out shortcuts. It also shows competitive performance compared to GPT-4o LMExaminer which has the highest correlation. All other baselines perform worse than GEM, and only BARTScore-Recall has a significant positive correlation. In later sections, we provide further comparisons between GEM and LMExaminer.
4.2 Result 2: Better Sensitivity to Degradation
We now compare how sensitive the evaluation metrics are to degradations that obviously reduce the semantic informativeness of the responses. A statistically significant score decrease indicates the sensitivity of the metric. We first introduce our validation workflow in Algorithm 2, the ICLR peer review dataset, and then introduce several degradation strategies. The workflow and dataset will also be used in the robustness check again manipulation strategies in Section 4.3.
@@ 4.2 Result 2: Better Sensitivity to Degradation
ICLR Dataset. We use the ICLR 2023333At the time of our experiment, ICLR 2024 review data was not yet available through the OpenReview API. peer review data publicly available on OpenReview. We randomly select 300 papers as our dataset, and for each paper, we randomly select 3 original human reviews: one review to serve as a human candidate, and the other two as reference responses.
Degradation Strategies. We introduce three degradation strategies as follows. Note that the degraded responses after Sentence Deletion or Deletion & Completion can be regarded as Blackwell-dominated by the original one. The detailed setup is provided in Appendix D.3.
  1. 1.
@@
Statistics and Results. We compare the standardized mean differences (SMD) of evaluation metrics in Table 2: GEM and GEM-S are only two metrics that consistently have significant negative SMD on all three degradations. The LMExaminer shows the highest effectiveness on Sentence Deletion and Deletion & Completion, but failed to penalize reviews only based on the abstract.
@@
Table 2: Standardized Mean Differences (SMD) $d:=(\mu^{\prime}-\mu)/{\sigma_{\text{pooled}}}$ of scores after degradations and manipulations of various evaluation metrics. Significant score decreases ($p<0.05$) after degradations are highlighted in bold green, implying the evaluation metric can effectively penalize the degradation. Significant score increases ($p<0.05$) after manipulations are highlighted in bold red, implying the evaluation metric is not robust against the manipulation. We provide 95% confidence interval of SMD values in Appendix A.1.
4.3 Result 3: Better Robustness against Manipulation